Kim's age in 6 years is triple her age 14 years ago. What is Kim's age?
Answer by macston(5194)

.
K=Kim's age
.
K+6=3(K-14)
K+6=3K-42
48=2K
K=24
ANSWER: Kim is 24 years old
.
CHECK:
K+6=3(K-14)
24+6=3(24-14)
30=3(10)
30=30